OpenStudy (anonymous):

simplify ^4√100 divided by ^4√5. ive read everything I could on how to solve it and I cant figure it out please help. the 4 are superscript and about the randicand

OpenStudy (ipwnbunnies):

\[\frac{\sqrt[4]{100}}{\sqrt[4]{5}} \] = \[\sqrt[4]{\frac{100}{5}} = \sqrt[4]{20}\]

OpenStudy (anonymous):

I have 4 choices if that will help

OpenStudy (ipwnbunnies):

Can't you figure it out from that? The fourth root of 20. You can't really break it down anymore.

OpenStudy (anonymous):

A. 5 ^4√2 B. 2 ^4√5 C, 2√5 D.^4√80

OpenStudy (anonymous):

those are the options it gives me

OpenStudy (ipwnbunnies):

Wat. None of these are right. ;-; Did you write the problem correctly?

OpenStudy (anonymous):

yes it is exponent 4√100 divided by exponent 4√5

OpenStudy (ipwnbunnies):

There's no way. The fourth root of 20 is the farthest you can go, with rational number. :/

OpenStudy (anonymous):

my answer was 16√2 but Idk if I can simplify it more??

OpenStudy (ipwnbunnies):

That'd still be incorrect. Idk dude. Broken question. :3

OpenStudy (anonymous):

or could it be D because it cant be broken down

OpenStudy (ipwnbunnies):

But 100/5 is 20. ._.

OpenStudy (anonymous):

can you teach my how to simplify A so that I can try and figure out the answer that way??

OpenStudy (ipwnbunnies):

Idk what you mean. .-. I did all the work in my first post.

OpenStudy (anonymous):

no I mean if you look at choice A how do you figure that out and find out how you get to that point

OpenStudy (johnweldon1993):

As long as your question looks exactly as @iPwnBunnies has shown on the first post of the thread...then there is no provided answer choice that is correct.

OpenStudy (anonymous):

when I talked to my teacher she said there was and she wasn't a big help

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

ok I just checked my email and she said that the 100 is a 400

OpenStudy (anonymous):

dooh!!!!

OpenStudy (anonymous):

yeah sorry she just said it was a typo and she apologized for confusing me lol

OpenStudy (neer2890):

\[4\sqrt{\frac{ 400 }{ 5 }}=4\sqrt{80}\]

OpenStudy (neer2890):

please consider that 4 is exponent here.

OpenStudy (anonymous):

hmmm.... B and D have the same value but B is simplified.... ie, \(\large 2\sqrt[4]5=\sqrt[4]{80} \)

OpenStudy (anonymous):

the answer is D correct?

OpenStudy (anonymous):

D has the same VALUE as B.... (verify that on your calculator) B is the simplified answer... if I had to choose, i'd choose B

OpenStudy (neer2890):

yup.. @dpaInc is correct..!

OpenStudy (anonymous):

so B is correct can you show me how B is correct i can get to 4^ square of 80 but how do i simplify??

OpenStudy (anonymous):

\(\large \sqrt[4]{80}=\sqrt[4]{16\cdot 5}=\sqrt[4]{16}\cdot \sqrt[4]{ 5}=2\cdot \sqrt[4]{ 5} \)

OpenStudy (anonymous):

how do you get the 2?

OpenStudy (anonymous):

\(\large \sqrt[4]{16}=2 \) because 2*2*2*2 = 16

OpenStudy (anonymous):

oh ok yeah i forgot thank you for helping me

OpenStudy (anonymous):

yw... :)